# On Springer correspondence for wreath products

### v1

@@ Abstract.
We first show that the wreath product $\Sigma_{m}\wr\Sigma_{d}$ between two symmetric groups appears as the generalized Weyl group of an Iwahori’s generalized Tits system. We then introduce a certain subvariety of the flag variety of type A, and then give a geometric proof of its Bruhat decomposition indexed by $\Sigma_{m}\wr\Sigma_{d}$, via the Bialynicki-Birula decomposition. Furthermore, we realize the group algebra $\mathbb{Q}[\Sigma_{m}\wr\Sigma_{d}]$ as the top Borel-Moore homology of a Steinberg variety. Such a geometric realization leads to a Springer correspondence for the irreducible representations over $\mathbb{C}[\Sigma_{m}\wr\Sigma_{d}]$, which can be regarded as a counterpart of the Clifford theory for wreath products.
@@ 1.1. Wreath Products between Symmetric Groups
On one hand, it is well-known that the Ariki-Koike algebras are cyclotomic quotients of the affine Hecke algebras of type A. It is standard (see [CG97]) that the affine Hecke algebras of type A can be obtained via the equivariant K-theory on the Steinberg varities. On the other hand, a classical (i.e., $q\mapsto 1$) result is also available by considering the Borel-Moore homology.
@@ 1.4. Main results
    (Corollary 2.3.4, Theorem 3.3.3) There is a Bruhat decomposition of $G$ with Bruhat cells indexed by $\Sigma_{m}\wr\Sigma_{d}$, for some algebraic groups $B,G$. Moreover, it affords a geometric variant $G/B=\bigsqcup_{w\in\Sigma_{m}\wr\Sigma_{d}}BwB/B$ via the Bialynicki-Birula decomposition.
@@
We plan to pursue the (equivariant) K-theory of our Steinberg variety $Z_{m\wr d}$ and its connection with the Hu algebras in a sequel.
@@ 4.1. Nilpotent Orbits
4.2. A Variant of Springer Resolution
@@ 6. Irreducible representations
6.1. Ginzburg’s Lagrangian Construction
We aim to apply the classification theorem [CG97, Theorem 3.5.7] in Chriss-Ginzburg. First, we recall the required assumptions listed in [CG97, §3.5]:
@@ 6.2. A Springer-type Classification of Irreducibles
Theorem 6.2.3.
The complete set of isomorphism classes of simple $\mathbb{C}[\Sigma_{m}\wr\Sigma_{d}]$-modules is given by $\{H^{\mathrm{BM}}_{\mathrm{top}}(\mathfrak{B}_{x})_{\psi}~{}|~{}[x,\psi]\in I^{G}_{m\wr d}\}$, where $I^{G}_{m\wr d}:=\{G_{m\wr d}\textup{-conjugacy class }[x,\psi]~{}|~{}x\in\mathcal{N}^{d}_{m},\ \psi\in C(x)^{\wedge})\}$.
@@ 6.3. Clifford Theory
$S^{\lambda}:=\bigotimes_{\nu\vdash m}S^{\lambda(\nu)}\in\mathop{\mathrm{Irr}}\mathbb{C}[\Sigma_{\gamma}],\quad S^{\gamma}:=\bigotimes_{\nu\vdash m}(S^{\nu})^{\otimes\gamma(\nu)}\in\mathop{\mathrm{Irr}}\mathbb{C}[\Sigma_{m}^{d}].$ (6.3.2)
@@
Proposition 6.3.1.
@@ 6.4. Characterization of the Index Set
In this subsection, we give an explicit characterization for the index set of the simple $\mathbb{C}[\Sigma_{m}\wr\Sigma_{d}]$-modules from the geometric approach in Theorem 6.2.3 and provide an identification to the index set in the algebraic approach from the above subsection.
From Theorem 6.2.3, each irreducible is indexed by a $G_{m\wr d}$-conjugacy class $[x,\psi]$ for some $x=(n_{i})_{i}\in\mathcal{N}_{m}^{d}$ and $\psi\in C(x)^{\wedge}$. Each $n_{i}\in\mathcal{N}_{m}$ has Jordan type $x_{i}\vdash m$. Equivalently, $x$ is associated with a map $\gamma(x)\in J_{m\wr d}$ given by
@@
As a consequence, we can now identify the irreducibles arising from Clifford theory with those arising from Springer theory.
@@
There assignment $\lambda\mapsto[x(\lambda),S^{\lambda}]$ defines a bijection $\Psi:I^{C}_{m\wr d}\to I^{G}_{m\wr d}$, where $x(\lambda)$ is a fixed element in $\mathcal{N}_{m}^{d}$ such that $\gamma(x(\lambda))=|\lambda|$.
Conversely, for any $[x,\psi]\in I^{G}_{m\wr d}$, $\psi\in\mathop{\mathrm{Irr}}\mathbb{C}[\Sigma_{\gamma(x)}]$ is of the form $S^{\lambda}:=\bigotimes_{\nu\vdash m}S^{\lambda(\nu)}$ for some $\lambda\in I^{C}_{m\wr d}$. Then, the inverse map is given by $[x,\psi]\mapsto\lambda$.
6.5. Special Cases
First, we consider an extreme case that $m=2$ so that the wreath product $\Sigma_{2}\wr\Sigma_{d}$ is a Weyl group of type $B_{d}$.
Example 6.5.1.
The index set $I^{C}_{2\wr d}=\{\lambda:\{(2),(1,1)\}\to\Pi~{}|~{}|\lambda((2))|+|\lambda((1,1))|=d\}$ is in bijection with the set of all bipartitions of $d$ via $\lambda\mapsto(\lambda((2)),\lambda((1,1)))$.
For the geometric index set $I^{G}_{2\wr d}$, note that any $x\in\mathcal{N}_{2}^{d}$ has Jordan type $(2)^{a}(1,1)^{d-a}$ for some $a$. The component group $C_{G_{2\wr d}}(x)$ is then isomorphic to the Young subgroup $\Sigma_{a}\times\Sigma_{d-a}$, and hence any $\psi\in C(x)^{\wedge}$ is of the form $S^{\lambda^{\prime}}\otimes S^{\lambda^{\prime\prime}}$ for some $\lambda^{\prime}\vdash a,\lambda^{\prime\prime}\vdash d-a$. Therefore, $\Psi$ sends $\lambda$ to $[x,S^{\lambda((2))}\otimes S^{\lambda((1,1))}]$.
Next, we consider the other extreme case when $d=2$. The wreath product $\Sigma_{m}\wr\Sigma_{2}$ is generally not a Coxeter group. We remark that irreducibles for the Hu algebras share the same index set $I^{H}_{m\wr 2}:=\{[\nu^{\prime},\nu^{\prime\prime}]~{}|~{}\nu^{\prime}\neq\nu^{\prime\prime}\in\Pi_{m}\}\sqcup\{[\nu,\nu]_{+},[\nu,\nu]_{-}~{}|~{}\nu\in\Pi_{m}\}$ with the irreducibles for $\mathbb{C}[\Sigma_{m}\wr\Sigma_{2}]$ (See [Hu02, LNX23]), where $[\nu^{\prime},\nu^{\prime\prime}]$ denotes the equivalence class in $\Pi_{m}^{2}$ under the obvious $\Sigma_{2}$-action.
Example 6.5.2.
The index set $I^{C}_{m\wr 2}=\{\lambda:\Pi_{m}\to\Pi~{}|~{}\sum_{\nu\vdash m}|\lambda(\nu)|=2\}$ is in bijection with $I^{H}_{m\wr 2}$ via
@@ 6.5. Special Cases
On the other hand, any $x\in\mathcal{N}_{m}^{2}$ either has Jordan type $\nu^{2}$ or $\nu^{\prime}\nu^{\prime\prime}$ for some $\nu\in\Pi_{m}$ or $\nu^{\prime}\neq\nu^{\prime\prime}\in\Pi_{m}$. For the former case, $C(x)\cong\Sigma_{2}$ and hence $x$ corresponds to two distinct elements $[x,S^{(2)}]$ and $[x,S^{(1,1)}]$ in $I^{G}_{m\wr 2}$. For the latter case, $C(x)\cong\Sigma_{1}\times\Sigma_{1}$ and hence $x$ corresponds exactly one element $[x,S^{(1)}]$ in $I^{G}_{m\wr 2}$.
Finally, $\Psi$ sends $\lambda$ to $[x,S^{\lambda(\nu)}]$ if $|\lambda(\nu)|=2$ for some $\nu\vdash m$, $x\in\mathcal{N}_{m}^{2}$ with Jordan type $\nu^{2}$; and it sends $\lambda$ to $[x,S^{(1)}]$ if $\lambda(\nu^{\prime})=\lambda(\nu^{\prime\prime})=(1)$ for some $\nu^{\prime}\neq\nu^{\prime\prime}\vdash m$, $x\in\mathcal{N}_{m}^{2}$ with Jordan type $\nu^{\prime}\nu^{\prime\prime}$.

### v2

# On the Springer correspondence for wreath products

@@ Abstract.
We first show that the wreath product $\Sigma_{m}\wr\Sigma_{d}$ between two symmetric groups appears as the generalized Weyl group of an Iwahori’s generalized Tits system. We then introduce a certain subvariety of the flag variety of type A, and then give a geometric proof of its Bruhat decomposition indexed by $\Sigma_{m}\wr\Sigma_{d}$, via the Bialynicki-Birula decomposition. Furthermore, we realize the group algebra $\mathbb{Q}[\Sigma_{m}\wr\Sigma_{d}]$ as the top Borel-Moore homology of a Steinberg variety. Such a geometric realization leads to a Springer correspondence for the irreducible representations over $\mathbb{C}[\Sigma_{m}\wr\Sigma_{d}]$, which can be regarded as a counterpart of the Clifford theory for wreath products. Consequently, we have obtained a new Springer correspondence of type B/C/D using essentially type A geometry.
@@ 1.1. Wreath Products between Symmetric Groups
1.2. The Lagrangian Construction
On one hand, it is well-known that the Ariki-Koike algebras are cyclotomic quotients of the affine Hecke algebras of type A, which can be obtained via the equivariant K-theory on the Steinberg varities (see [CG97, KL87]). On the other hand, a classical (i.e., $q\mapsto 1$) result is also available via a Lagrangian construction due to Ginzurg [Gi86] and Kashiwara-Tanisaki [KT84] by considering the Borel-Moore homology.
In [CG97], it is established a uniform geometric approach to the representation theory of the Weyl groups, their corresponding Lie algebras, and various quantizations of these objects. Such a Lagrangian construction appeared also in [Li21] for certian symmetric pairs. In this paper, we show that the wreath product $\Sigma_{m}\wr\Sigma_{d}$ is a new example on which the Lagrangian construction applies.
@@ 1.5. Main results
    (Corollary 2.3.4, Theorem 3.3.3) There is a Bruhat decomposition of $G$ with Bruhat cells indexed by $\Sigma_{m}\wr\Sigma_{d}$, for some algebraic groups $B,G$. Moreover, it affords a geometric variant $G/B=\bigsqcup_{w\in\Sigma_{m}\wr\Sigma_{d}}BwB/B$ via the Bialynicki-Birula decomposition [BB73].
@@
The Springer correspondence here is a generalization of Ginzburg’s construction [Gi86] (see also equivalent constructions of Springer [Sp76, Sp78], and of Lusztig [Lu81]). Finally, we remark that the obvious convolution algebras obtained by the wreath flag varieties over finite fields realizes the “Hecke algebra” in [I65] that is a twisted tensor product $\mathcal{H}_{q}(\Sigma_{m}^{d})\hat{\otimes}\mathbb{C}[\Sigma_{d}]$, instead of the Hu algebra. We plan to pursue the (equivariant) K-theory of our Steinberg variety $Z_{m\wr d}$ and its potential connection with the Hu algebras in a sequel.
@@ 4.1. Nilpotent Orbits
4.2. A Variant of the Springer Resolution
@@ 6. Irreducible representations
6.1. A Lagrangian Construction
We aim to apply the classification theorem (see [CG97, Theorem 3.5.7]). First, we recall the required assumptions listed in [CG97, §3.5]:
@@ 6.2. A Springer-type Classification of Irreducibles
Theorem 6.2.3 (Springer correspondence for $\Sigma_{m}\wr\Sigma_{d}$).
The complete set of isomorphism classes of simple $\mathbb{C}[\Sigma_{m}\wr\Sigma_{d}]$-modules is given by $\{H^{\mathrm{BM}}_{\mathrm{top}}(\mathfrak{B}_{x})_{\psi}~{}|~{}[x,\psi]\in I^{S}_{m\wr d}\}$, where
$I^{S}_{m\wr d}:=\{G_{m\wr d}\textup{-conjugacy class }[x,\psi]~{}|~{}x\in\mathcal{N}^{d}_{m},\ \psi\in C(x)^{\wedge})\}.$ (6.2.6)
@@ 6.3. Clifford Theory
$S^{\lambda}:={\textstyle\bigotimes_{\nu\vdash m}}S^{\lambda(\nu)}\in\mathop{\mathrm{Irr}}\mathbb{C}[\Sigma_{\gamma}],\quad S^{\gamma}:={\textstyle\bigotimes_{\nu\vdash m}}(S^{\nu})^{\otimes\gamma(\nu)}\in\mathop{\mathrm{Irr}}\mathbb{C}[\Sigma_{m}^{d}].$ (6.3.2)
@@
Proposition 6.3.1 (Clifford theory for $\Sigma_{m}\wr\Sigma_{d}$).
@@ 6.4. Characterization of the Index Set
For completeness, in this subsection, we give an explicit characterization for the index set $I^{S}_{m\wr d}$ from the Springer correspondence (Theorem 6.2.3), and then provide an identification with the index set $I^{C}_{m\wr d}$ in the algebraic approach from Clifford theory.
From (6.2.6), each irreducible is indexed by a $G_{m\wr d}$-conjugacy class $[x,\psi]$ for some $x=(n_{i})_{i}\in\mathcal{N}_{m}^{d}$ and $\psi\in C(x)^{\wedge}$. Each $n_{i}\in\mathcal{N}_{m}$ has Jordan type $x_{i}\vdash m$. Equivalently, $x$ is associated with a map $\gamma(x)\in J_{m\wr d}$ given by
@@
As a consequence, we can now identify the irreducibles arising from Clifford theory with those arising from Springer theory. Recall the sets $I^{S}_{m\wr d}$, $I^{C}_{m\wr d}$, and the module $S^{\lambda}$ from (6.2.6), (6.3.1), and (6.3.2), respectively.
@@
There is a natural bijection $\Psi:I^{C}_{m\wr d}\to I^{S}_{m\wr d}$ between the index sets. $\Psi$ and its inverse map are given, respectively, by
$\Psi(\lambda)=[x_{\lambda},S^{\lambda}],\quad\Psi^{-1}([x,\psi])=\lambda_{\psi},$ (6.4.9)
where $x_{\lambda}\in\mathcal{N}_{m}^{d}$ is an element of Jordan type $(x_{i})_{i}\in\Pi_{m}^{d}$ such that $|\lambda(\nu)|=\gamma(x_{\lambda}):=\#\{1\leq i\leq d~{}|~{}x_{i}=\nu\}$ for all $\nu\vdash m$, and $\lambda_{\psi}$ is the element in $I^{C}_{m\wr d}$ such that $\psi\cong S^{\lambda_{\psi}}:=\bigotimes_{\nu\vdash m}S^{\lambda_{\psi}(\nu)}\in\mathop{\mathrm{Irr}}\mathbb{C}[\Sigma_{\gamma(x)}]$.
Proof.
It follows from the construction that $\Psi$ is well-defined. On the other hand, any $\psi\in C(x)^{\wedge}$ is isomorphic to $S^{\lambda_{\psi}}$ thanks to Proposition 6.4.2. It is routine to check that $\Psi$ and $\Psi^{-1}$ are inverse to each other. ∎
6.5. Extreme Cases
First, we consider an extreme case that $m=2$ so that the wreath product $\Sigma_{2}\wr\Sigma_{d}$ is a Weyl group of type $B_{d}=C_{d}$. Our theorem (that uses a essentially type A geometry) gives a new Springer correspondence for Weyl groups of type B/C.
Example 6.5.1 (new Springer correspondence for type B/C).
Let $\Pi_{d}^{(2)}:=\{(\nu^{\prime},\nu^{\prime\prime})~{}|~{}\nu^{\prime}\vdash a,\nu^{\prime\prime}\vdash d-a\}$ be the set of bipartitions of $d$. The index set $I^{C}_{2\wr d}=\{\lambda:\{(2),(1,1)\}\to\Pi~{}|~{}|\lambda((2))|+|\lambda((1,1))|=d\}$ is in bijection with $\Pi_{d}^{(2)}$ via $\lambda\mapsto(\lambda((2)),\lambda((1,1)))$.
For the geometric index set $I^{S}_{2\wr d}$, note that any $x\in\mathcal{N}_{2}^{d}$ has Jordan type $(2)^{a}(1,1)^{d-a}$ for some $a$. The component group $C_{G_{2\wr d}}(x)$ is then isomorphic to the Young subgroup $\Sigma_{a}\times\Sigma_{d-a}$, and hence any $\psi\in C(x)^{\wedge}$ is of the form $S^{\lambda^{\prime}}\otimes S^{\lambda^{\prime\prime}}$ for some $\lambda^{\prime}\vdash a$, $\lambda^{\prime\prime}\vdash d-a$. Therefore, we obtain the following bijections:
$\begin{array}[]{rcl}\mathop{\mathrm{Irr}}\mathbb{C}[W_{B_{d}}]\equiv\mathop{\mathrm{Irr}}\mathbb{C}[W_{C_{d}}]\equiv\Pi_{d}^{(2)}\equiv I^{C}_{2\wr d}&\to&I^{S}_{2\wr d},\\ \lambda&\mapsto&[x,S^{\lambda((2))}\otimes S^{\lambda((1,1))}].\end{array}$ (6.5.1)
Next, we consider the other extreme case when $d=2$. The wreath product $\Sigma_{m}\wr\Sigma_{2}$ is generally not a Coxeter group. We remark that irreducibles for the Hu algebras share the same index set
$I^{H}_{m\wr 2}:=\{[\nu^{\prime},\nu^{\prime\prime}]~{}|~{}\nu^{\prime}\neq\nu^{\prime\prime}\in\Pi_{m}\}\sqcup\{[\nu,\nu]_{+},[\nu,\nu]_{-}~{}|~{}\nu\in\Pi_{m}\}$ (6.5.2)
with the irreducibles for $\mathbb{C}[\Sigma_{m}\wr\Sigma_{2}]$ (See [Hu02, LNX23]), where $[\nu^{\prime},\nu^{\prime\prime}]$ denotes the equivalence class in $\Pi_{m}^{2}$ under the obvious $\Sigma_{2}$-action. Note that each equivalence class of the form $[\nu,\nu]$ corresponds to two simple modules, and hence one can index these modules (called $D(\nu)_{+}$ and $D(\nu)_{-}$ in [Hu02, §4]) by $[\nu,\nu]_{+}$ and $[\nu,\nu]_{-}$, respectively.
Example 6.5.2 (The wreath product $\Sigma_{m}\wr\Sigma_{2}$).
Recall that $I^{C}_{m\wr 2}=\{\lambda:\Pi_{m}\to\Pi~{}|~{}\sum_{\nu\vdash m}|\lambda(\nu)|=2\}$. We have a bijection $I^{C}_{m\wr 2}\to I^{H}_{m\wr 2}$ given by
@@ 6.5. Extreme Cases
On the other hand, any $x\in\mathcal{N}_{m}^{2}$ either has Jordan type $\nu^{2}$ or $\nu^{\prime}\nu^{\prime\prime}$ for some $\nu\in\Pi_{m}$ or some distinct partitions $\nu^{\prime}\neq\nu^{\prime\prime}\in\Pi_{m}$. For the former case, $C(x)\cong\Sigma_{2}$ and hence $x$ corresponds to two distinct elements $[x,S^{(2)}]$ and $[x,S^{(1,1)}]$ in $I^{S}_{m\wr 2}$. Note that the plus and minus signs in (6.5.2) now correspond to a choice of the trivial module $S^{(2)}$ or the sign module $S^{(1,1)}$ over $\mathbb{C}[\Sigma_{2}]$. For the latter case, $C(x)\cong\Sigma_{1}\times\Sigma_{1}$ and hence $x$ corresponds to exactly one element $[x,S^{(1)}]$ in $I^{S}_{m\wr 2}$.
Finally, $\Psi$ sends $\lambda$ to $[x,S^{\lambda(\nu)}]$ if $|\lambda(\nu)|=2$ for some $\nu\vdash m$ and some $x\in\mathcal{N}_{m}^{2}$ with Jordan type $\nu^{2}$; and it sends $\lambda$ to $[x,S^{(1)}]$ if $\lambda(\nu^{\prime})=\lambda(\nu^{\prime\prime})=(1)$ for some $\nu^{\prime}\neq\nu^{\prime\prime}\vdash m$ and some $x\in\mathcal{N}_{m}^{2}$ with Jordan type $\nu^{\prime}\nu^{\prime\prime}$.
Remark 6.5.3 (new Springer correspondence for type D).
Combining Example 6.5.2 with the Morita equivalence result, the index set of the irreducibles for the Weyl group (and hence the Hecke algebra) of type $D$ are in bijection with the following set:
$I^{H}_{D_{d}}:=\begin{cases}\{[\nu^{\prime},\nu^{\prime\prime}]~{}|~{}(\nu^{\prime},\nu^{\prime\prime})\in\Pi_{d}^{(2)},\ \nu^{\prime}\neq\nu^{\prime\prime}\}&\textup{if }d=2m+1;\\ \{[\nu^{\prime},\nu^{\prime\prime}]~{}|~{}(\nu^{\prime},\nu^{\prime\prime})\in\Pi_{d}^{(2)},\ \nu^{\prime}\neq\nu^{\prime\prime}\}\sqcup\{[\nu,\nu]_{+},[\nu,\nu]_{-}~{}|~{}(\nu,\nu)\in\Pi_{d}^{(2)}\}&\textup{if }d=2m.\end{cases}$ (6.5.4)
Once again, we obtain a new Springer correspondence for type D using geometry of type A. Let $\mathcal{N}_{d}^{(2)}:=\bigsqcup_{1\leq a\leq d}\mathcal{N}_{a}\times\mathcal{N}_{d-a}$. For $x\in\mathcal{N}_{d}^{(2)}$, let $(\nu^{\prime}_{x},\nu^{\prime\prime}_{x})\in\Pi_{d}^{(2)}$ be the corresponding Jordan type. Define
$I^{S}_{D_{d}}:=\begin{cases}\{[x,S^{(1)}]~{}|~{}x\in\mathcal{N}_{d}^{(2)},\nu^{\prime}_{x}\neq\nu^{\prime\prime}_{x}\}&\textup{if }d=2m+1;\\ \{[x,S^{(1)}]~{}|~{}x\in\mathcal{N}_{d}^{(2)},\nu^{\prime}_{x}\neq\nu^{\prime\prime}_{x}\}\sqcup\{[x,S^{(2)}],[x,S^{(1,1)}]~{}|~{}x\in\mathcal{N}_{d}^{(2)},\nu^{\prime}_{x}=\nu^{\prime\prime}_{x}\}&\textup{if }d=2m.\end{cases}$ (6.5.5)
Therefore, we obtain a bijection $\mathop{\mathrm{Irr}}\mathbb{C}[W_{D_{d}}]\equiv I^{H}_{D_{d}}\to I^{S}_{D_{d}}$ given by
$[\nu^{\prime},\nu^{\prime\prime}]\mapsto[x,S^{(1)}],\quad[\nu^{\prime},\nu^{\prime\prime}]_{\epsilon}\mapsto\begin{cases}{}[x,S^{(2)}]&\textup{if }\nu^{\prime}=\nu^{\prime\prime},\ \epsilon=+;\\ {}[x,S^{(1,1)}]&\textup{if }\nu^{\prime}=\nu^{\prime\prime},\ \epsilon=-,\end{cases}$ (6.5.6)
where $x\in\mathcal{N}^{(2)}_{d}$ is some element of Jordan type $\nu^{\prime}\nu^{\prime\prime}$ with $(\nu^{\prime},\nu^{\prime\prime})\in\Pi^{(2)}_{d}$.
Finally, we remark that the original Springer correspondence can be described via the Lusztig-Shoji algorithm [Lu79, Sh79, Sh83] in terms of type D geometry. See also [SW19] for a diagrammatic approach to Springer correspondence for type D, for two-row partitions.